I need help with this one. I came up with three different answers!
-7/8x = -9/10

Answer by rfer(16322)

-7/8x=-9/10
x=(-9/10)/(-7/8)
invert and multiply
x=-9/10*-8/7
x=72/70
x=1 1/35